OpenStudy (anonymous):

f(x)=I x-1I How to find the x-intercept and y-intercept. Pls help me

OpenStudy (michele_laino):

step#1 we have to set x=0, what do you get?

OpenStudy (anonymous):

y= \[\left| -1 \right|\] y= 1

OpenStudy (michele_laino):

sorry, we have to use another method

OpenStudy (anonymous):

okay so what do we do?

OpenStudy (anonymous):

are you trying to graph it?

OpenStudy (michele_laino):

I can rewrite your function as below: \[\begin{gathered} f\left( x \right) = x - 1,\quad x - 1 \geqslant 0 \hfill \\ f\left( x \right) = 1 - x,\quad x - 1 < 0 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

yes @Ghuman499

OpenStudy (michele_laino):

since I have applied the definition of absolute value of a number

OpenStudy (anonymous):